[Brahms-l] Shift report 20050509 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon May 09 2005 - 08:06:09 EDT
Shifter leader was SB
Shifters were 


1. First new store at 2:00 am aborted about 15 minutes later due to RHIC Quench link problem. Second new store around 4:50 am and still on.

2. FS @2.3 deg @3450A , MRS @90 deg @1050A

trig2(424k), trig3(154k), trig4(607k), trig5(602k), trig6(500k)

3. No detector problems
-----------------------------------------------
Time(AM)     Blue Pol.(%)   Yellow Pol.(%)
-----------------------------------------------
2:00          48.8+/-2.9      46.6+/-2.7
5:00          49.2+/-2.9      45.6+/-2.7
6:50          50.7+/-2.9      49.4+/-2.6
-----------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon May 9 08:06:25 2005

This archive was generated by hypermail 2.1.8 : Mon May 09 2005 - 08:06:37 EDT